OpenStudy (anonymous):

Question: 13 of 27: Simplify the expression 3 + 9 × 4 ÷ 2. Select one of the options below as your answer: * A. 24 * B. 21 * C. 16 * D. 19

OpenStudy (anonymous):

thanks

myininaya (myininaya):

3+9*4/2 =3+36/2 =3+18 =21

myininaya (myininaya):

its B

OpenStudy (anonymous):

^^^ correct. sorry i can't add.

OpenStudy (anonymous):

darn already put it in

OpenStudy (anonymous):

i'm sorry! :( serious brain fart right there. sorry you had to suffer the consequences.